graph of y = x³ + 6 is translated 4 units to the right.



The translated graph has equation y = f(x)



Work out f(x).



Give your answer in the form x³+ ax² + bx + c where a, b and c are integers.

1
Expert's answer
2021-12-30T06:33:31-0500

Translation of graph means to increase the graph in respected axis


It may be x -Axis or y-Axis .




So, according to the. Question

Y= X3+6 is translated 4 units to the right.



Y=(x-4)+6

Identity of (x-a)3 = x-a3-3x2a +3xa2

Y=(x3-64-(3x2*4)+3x*16 ) +6

Y= x3-12x+48x -58



So, the value of

a is - 12

b is 48

c is -58